MICROWAVE STUDIES OF COMPLEXES OF METHANOL WITH CARBON DIOXIDE AND FORMALDE-HYDE

\underline{F. J. LOVAS}, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441; V. V. ILYUSHIN, Institute of Radio Astronomy of NASU, Chervonopraporna 4, 61002 Kharkov, Ukraine; and D. F. PLUSQUELLIC, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

The dimeric complexes, CH_{3}OH-CO_{2} and CH_{3}OH-H_{2}CO, were observed with pulsed-beam Fourier transform spectrometers at NIST. For the formaldehyde complex a heated reservoir nozzle was emploved and pressurized with a gas sample of 1% methanol in either Ne or Ar carrier and the nozzle heated to about 75 ^{∘}C. For the CO_{2} complex, a cylinder containing 1% methanol and 1% CO_{2} in Ar was used with a room temperature nozzle. Measurements were carried out from 8 GHz to 23 GHz. Both complexes exhibited internal rotation splitting due to internal rotation of the methyl group, and the formaldehyde complex showed an additional splitting on b-type transitions which is interpreted as a C_{2} tunneling motion of the H_{2}CO monomer. Several isotopically substituted forms have also been studied for improved structure determination. The CO_{2} complex is van der Waals bonded (C of carbon dioxide to O of the methanol), while the formaldehyde complex is hydrogen bonded (H of the hydroxyl group to O of formaldehyde). Discussion of the rotational, Stark, internal rotation, and structure analyses will be presented. These results will be compared to ab initio calculations, as well.